OpenStudy (anonymous):

Which of the following statements is FALSE? 1. The total amount of energy and matter in the Universe is constant.

OpenStudy (anonymous):

E=mc^2

OpenStudy (anonymous):

That has so got to be false. With exploding stars and asteroids striking other objects in space. The amount of energy has got to change at some point, and probably most of the time.

OpenStudy (anonymous):

E=mc^2

OpenStudy (anonymous):

So energy can be constant

OpenStudy (kainui):

I mean, the total amount of energy and matter in the universe is constant seems to me to be neither something a school kid, or anyone for that matter, be able to say fairly surely that there is a true or false answer to. Seriously.

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

This is just a back and fourth debatable question, if you put it into google, you'll see 1000s of opinions on this.

OpenStudy (anonymous):

Conventionally, we believe energy is conserved. We believe matter is conserved, except when transformed into energy through E=mc^2 processes. Thus, matter + enegy = conserved quantity.